# An Event Structure-aware Generative Model for Biomedical Event Extraction

### v3

@@ Abstract
Biomedical Event Extraction (BEE) is a challenging task that involves modeling complex relationships between fine-grained entities in biomedical text. BEE has traditionally been formulated as a classification problem. With the recent technological advancements in large language models (LLMs), generation-based models that cast event extraction as a sequence generation problem have attracted much attention from the NLP research communities. However, current generative models often overlook the importance of cross-instance information from complex event structures such as nested events and overlapping events, which contribute quite significantly in the benchmark datasets. In this paper, we propose an event structure-aware generative model called GenBEE, which can capture complex event structures in biomedical text for biomedical event extraction. In particular, GenBEE constructs event prompts that distill knowledge from LLMs for incorporating both label semantics and argument dependency relationships into the proposed model. In addition, GenBEE also generates prefixes with event structural prompts to incorporate structural features for improving the model’s overall performance. We have evaluated the proposed GenBEE model on three widely used biomedical event extraction benchmark datasets, namely MLEE, GE11, and PHEE. Experimental results show that GenBEE has achieved state-of-the-art performance on the MLEE and GE11 datasets, and achieved competitive results when compared to the state-of-the-art classification-based models on the PHEE dataset.
Index Terms:
@@
Biomedical Event Extraction, Generative Models, Structural Information, Large Language Models, Knowledge Distillation
I Introduction
@@
Biomedical Event Extraction (BEE) is a challenging task that involves identifying molecular events from natural language text, where an event typically includes certain biomedical entities, such as genes, proteins, or cellular components [1]. BEE is typically divided into two subtasks, namely biomedical event trigger detection and biomedical event argument extraction. Biomedical event trigger detection aims to identify trigger words and classify them by event types that represent the presence of biomedical events. On the other hand, biomedical event argument extraction aims to identify the arguments and associate them with specific roles in biomedical events, correlating them with the corresponding event triggers. BEE can provide invaluable information to facilitate the curation of knowledge bases and biomolecular pathways [2].
Traditionally, BEE has been formulated as a classification problem [3, 4]. However, these classification-based methods often struggle with encoding label semantics and other forms of weak supervision information [5, 6, 7]. Many current methods extract events in a pipeline manner, where trigger detection and argument extraction are treated as separate phases, which hinders the learning of shared knowledge between subtasks [8]. Recently, generation-based models have been proposed for event extraction (EE) on general domains [9, 10]. These models typically cast the event extraction task as a sequence generation problem using encoder-decoder based pre-trained language models (PLMs) [11, 12] to output conditional generation sequences. These generation-based models benefit from natural language prompts to learn additional semantic information [13]. Apart from employing natural language prompts to enhance semantic learning from the given context, some recent works such as GTEE [14] and AMPERE [15] have also proposed injecting continuous prompts [16, 17] into their frameworks for better contextual representation learning for event argument extraction.
@@
However, most current generation-based EE models [14, 13, 15] have difficulties when tackling events with complex structures, such as nested events. Compared to flat event structures, texts in the biomedical domain often contain more nested event structures [18]. For example, in benchmark datasets like MLEE [19] and GE11 [20], the nested event structures constitute approximately 25% and 28% of all events in the datasets, respectively. Figure 1 shows an example of two nested events identified from the given text. The trigger word “promoters” of the Positive Regulation event (E2) serves as the argument role Site for the trigger word “bind”, which triggers the Binding event (E1). Furthermore, “TCF-1 alpha” serves as the argument role Theme for both events (E1 and E2). In addition, a trigger word may also trigger different events in a sentence, depending on its event type or arguments. Events with such structures are called overlapping events. Identifying these complex structures is important for BEE, as it can reveal the underlying relationships between events. For example, a Site2 argument in a Binding event is likely to be a trigger for a Positive Regulation event.
@@
For generation-based EE models, an intuitive way to incorporate structural information is through prompt engineering [21] by adding natural language prompts that describe complex structural information as part of the input to the generative language model’s encoder. This helps elicit the rich knowledge distributed in the language model. However, Fei et al. [22] pointed out that the mainstay contextual representations in a generative language model’s encoder tend to weaken other high-level information such as structural information, thereby reducing the effects of such information at the decoder. Moreover, it is also unclear whether the generative language model contains sufficient domain knowledge to represent complex biomedical events. Therefore, it is a challenging task to incorporate structural information effectively into generative BEE models.
@@
To address the above issues, we propose an event structure-aware Generative Biomedical Event Extraction model (GenBEE), which aims to capture complex event structures effectively from biomedical texts. First, GenBEE constructs type description prompts to encapsulate the semantics of type labels. Next, it distills the knowledge from large language models (LLMs), such as GPT-4 [23], by creating event template prompts, which incorporate information such as dependencies among argument roles, to enhance contextual representation learning. Additionally, GenBEE introduces a structural prefix learning module that generates structure-aware prefixes, enriched with structural prompts to infuse the structural features.
@@
GenBEE functions as a comprehensive BEE framework that handles event trigger detection and event argument extraction simultaneously, and effectively leverages the shared knowledge and dependencies between these two subtasks. The main contributions of this paper are summarized as follows:
@@
II Related Work
II-A Biomedical Event Extraction
Many current methods for biomedical event extraction adopt a “pipeline” strategy, where trigger detection and argument extraction are treated as separate phases. However, “pipeline” methods may result in propagation of errors. To address this issue, Trieu et al. [8] introduced an end-to-end method called DeepEventMine, which simultaneously identifies triggers and allocates roles to entities, thereby mitigating error propagation. However, DeepEventMine requires the full annotation of all entities, which may not always be accessible in datasets. Ramponi et al. [6] considered biomedical event extraction as a sequence labeling task that involves joint training models for trigger detection and argument extraction via multi-task learning. Aside from these classification-based methods, some recent works have formulated event extraction as a question-answering task [24, 25]. This new paradigm transforms traditional classification methods into multiple questioning rounds, producing a natural language answer about a trigger or an argument in each round. Current QA-based event extraction methods are primarily focused on formulating separate questions for different events and argument types. As such, these QA methods lack the capability on capturing dependency information between arguments, which could be effective for assigning argument types.
@@ II-A Biomedical Event Extraction
II-B Generative Event Extraction
Recently, another emerging line of work that treats event extraction as a conditional generation problem [26, 27, 28, 13] has evolved. The development of generation-based event extraction models begins with the reformulation of event extraction problems as a generation task. For example, TANL [26] treated event extraction as translation tasks with augmented labels by using a variety of brackets and vertical bar symbols for representation, and tuned the model to predict these augmented labels. TempGen [27] proposed using templates for constructing role-filler tasks for entity extraction, and generating outputs that fill role entities into non-natural templated sequences. To adapt generation-based models to the task of event argument extraction, BART-Gen [28] investigated generative document-level event argument extraction. However, these methods only use naive natural language prompts and have not considered the incorporation of semantic information into the natural language prompts.
To better incorporate label semantics into the model, DEGREE [13] proposed a template-based model that incorporates label semantics into the designed prompts to extract event information. More recently, researchers have started to advance this line of work. For example, GTEE [14] used prefixes [16], which are sequences of tunable continuous prompts, to incorporate contextual type information into the generation-based model. AMPERE [15] proposed to equip the generative model with AMR-aware prefixes to embed the AMR graph, thereby providing more contextual information. To the best of our knowledge, there is no work done on integrating event structural features of complex events into generative models for BEE. Therefore, in this paper, we propose a novel generative model called GenBEE that incorporates complex event structural information for BEE.
III Method
III-A Event Prompt Construction
TABLE I: Example event prompts for the GE11 dataset.
@@ III-A Event Prompt Construction
Phosphorylation Phosphorylation events capture the enzymatic process of attaching a phosphate group to a target protein. Event trigger {Trigger} <SEP> Enzyme at {ROLE_Site} catalyzes the phosphorylation of {ROLE_Theme}.
@@
Table I shows some example event prompts for the GE11 dataset. For example, the event type description for the Binding event is “Binding involves two or more molecules coming together to form a complex for various biological processes and signaling pathways.”. Note that the type description is obtained according to the textual description of each event type provided in the technical report of the corresponding biomedical event extraction dataset (e.g., GE11 [20]). In the event template, the trigger part is “Event trigger {Trigger}”, which remains the same for all event types. The argument part is specific to each event type $e_{i}$. Each {Role_<argument>} serves as a placeholder for the corresponding argument role for the event. For example, the argument part for the Binding event in the GE11 dataset is: {Role_Theme} at binding site {Role_Site} and {Role_Theme2} at adjacent site {Role_Site2} form a complex, assisted by {Role_Theme3} and {Role_Theme4}.”.
Most previous works [14, 15] proposed to use manually designed templates for event extraction. However, it is challenging and time-consuming to define event templates for each event type in the biomedical event datasets, as domain knowledge is required. In this work, we propose to employ the Large Language Model (LLM) GPT-4 [23] with designed prompts to generate the event templates for the event types pre-defined in the biomedical event datasets. Figure 3 shows an input prompt that we have designed for generating the event template for the Binding event using GPT-4. The input prompt consists of an “Instruction” that specifies the generation objective for the biomedical event extraction template and a “Basic Template” that outlines the target trigger and argument roles. The basic template is constructed based on the pre-defined roles given in the technical report associated with the biomedical event dataset. The output of GPT-4 is an updated event template for the corresponding event type, which contains richer information than the basic template for generative event extraction.
Figure 3: Prompts for eliciting knowledge from LLM to construct the event template for the Binding event in the GE11 dataset.
@@
III-B Structural Prefix Learning
In biomedical texts, events are often complex and interrelated. Identifying the latent relationships among entities across different events could provide important clues for capturing complex event structures, thereby enhancing representation learning. Table II shows the structural prompts [29] constructed in the proposed model to capture different event structures for the Binding event. The structural prompts aim to provide descriptions of different event structures and then elicit the PLM to identify complex event structures by capturing the latent relationships among entities across different biomedical events. As shown in Table II, the structural prompts are the same for all event types, except the corresponding event type name specified between the two tags of <T>. In this work, structural prompts are categorized into three groups: General Events, Overlapping Events, and Nested Events. The General Events describe the structural information of common events, guiding the PLM to identify potential interactions between frequently co-occurring events. The Overlapping Events describe the structural information of overlapping events, instructing the PLM to identify trigger words that commonly trigger overlapping events. The Nested Events describe the structural information of nested events, guiding the PLM to identify entities that frequently serve as argument roles in the events while acting as trigger words in others, or vice versa. Overall, the structural prompts are designed to describe the various structural information of biomedical events for generative event extraction.
TABLE II: Example structural prompts for the Binding event.
@@ III-B Structural Prefix Learning
General Events (S1) Explore events that frequently co-occur with <T>Binding<T> events, aiming to identify and analyze the interactions and dependencies among these events to enhance understanding of their interrelationships.
Overlapping Events (S2) Explore entities that serve as triggers in both <T>Binding<T> events and other event types, aiming to clarify the overlap in trigger roles across different contexts to better understand trigger versatility.
Nested Events (S3) Explore entities acting in multiple roles, including as roles in <T>Binding<T> events and differently in other events, highlighting the dynamics of role versatility and their implications for event structure.
(S4) Explore entities where the trigger of <T>Binding<T> events also acts as a role in other events, or vice versa, highlighting these complex inter-event relationships to identify patterns of event interaction.
@@
III-C Generative Model Training
@@
IV Experiments
TABLE III: Statistics of the datasets.
@@ IV Experiments
IV-A Datasets
We have conducted experiments on the following three publicly available benchmark datasets, namely MLEE, GE11, and PHEE, for biomedical event extraction. MLEE [19] has 29 event types and 14 argument roles. We use the train/dev/test split given by the data provider. GE11 [20] has 9 event types and 10 role arguments. We use the train/dev/test split given by the shared task and evaluate the performance based on the development set, as the test set is unannotated and the official tool for evaluation is no longer available. PHEE [30] has 2 event types and 16 role arguments. We use the train/dev/test split provided by TextEE [31] and preprocess each dataset accordingly. Table III presents the statistics of the datasets. For evaluation metrics, we follow the previous work [32, 10, 13] to adopt the following two metrics to evaluate the performance: (1) Trigger Classification (Trg-C) F1-score: A trigger is correctly classified if the predicted span and the predicted event type match the gold ones. (2) Argument Classification (Arg-C) F1 score: An argument is correctly classified if the predicted span, event type, and role type match the gold ones.
IV-B Implementation Details
For a fair comparison with recent works, we leverage the BART-large [12] as the PLM in our proposed GenBEE model. For encoding structural prompts, we use a pre-trained BioBERT [33] model. We train GenBEE on an NVIDIA A100 40G GPU. The learning rate of BART-large is set to $10^{-5}$ and the learning rate of BioBERT is set to $10^{-6}$. Moreover, we train GenBEE for 50 epochs on PHEE and 80 epochs on MLEE and GE11. The batch size is set to 16 during training. The prefix length $l$ is set to 40, which is determined experimentally within the interval $l$ = {20, 30, 40, 50, 60}.
IV-C Baseline Models
We compare our GenBEE with the following EE models: (1) DyGIE++ [32] is a classification-based model that captures contextual information with the span graph propagation technique. (2) OneIE [34] is a classification-based model that utilizes global feature-aware graphs to capture cross-subtask and cross-instance interactions. (3) AMR-IE [35] is a classification-based model that captures syntactic characteristics in contexts using the AMR graph. (4) EEQA [36] is a classification-based model that formulates EE as a question-answering task. (5) TagPrime [37] is a state-of-the-art classification-based EE model that iteratively extracts event information based on each event type. (6) DEGREE [13] is a state-of-the-art generation-based EE model that integrates prompts for conditional generation. As the vanilla DEGREE model was designed mainly for general domain event extraction, we have re-implemented it using the prompts provided by [31] for the MLEE, GE11, and PHEE datasets. To ensure a fair comparison across models, we adopt the official codes of the above baselines and train them with the same data.
TABLE IV: Experimental results (%) for extracting events based on the MLEE, GE11 and PHEE datasets. The best score is highlighted in bold, and the second-best score is underlined.

### v5

# A Structure-aware Generative Model for Biomedical Event Extraction

@@ Abstract
Biomedical Event Extraction (BEE) is a challenging task that involves modeling complex relationships between fine-grained entities in biomedical text. BEE has traditionally been formulated as a classification problem. With recent advancements in large language models (LLMs), generation-based models that cast event extraction as a sequence generation problem have attracted attention in the NLP research community. However, current generative models often overlook cross-instance information in complex event structures, such as nested and overlapping events, which constitute over 20% of events in benchmark datasets. In this paper, we propose GenBEE, an event structure-aware generative model that captures complex event structures in biomedical text for biomedical event extraction. GenBEE constructs event prompts that distill knowledge from LLMs to incorporate both label semantics and argument dependency relationships. In addition, GenBEE generates prefixes with event structural prompts to incorporate structural features to improve the model’s overall performance. We have evaluated the proposed GenBEE model on three widely used BEE benchmark datasets, namely MLEE, GE11, and PHEE. Experimental results show that GenBEE has achieved state-of-the-art performance on the MLEE and GE11 datasets, and achieved competitive results when compared to the state-of-the-art classification-based models on the PHEE dataset. Our source code is released at https://github.com/HaohanYuan01/GenBEE.
Keywords:
@@
Biomedical Event Extraction Generative Models Prefix Learning Large Language Models Knowledge Distillation
1 Introduction
@@
Biomedical Event Extraction (BEE) is a challenging task that involves identifying molecular events from natural language text, which can provide invaluable information to facilitate the curation of knowledge bases and biomolecular pathways [15]. BEE is typically divided into two subtasks, namely biomedical event trigger detection and biomedical event argument extraction. Biomedical event trigger detection aims to identify trigger words and classify them by event types that represent the presence of biomedical events. On the other hand, biomedical event argument extraction aims to identify the arguments and associate them with specific roles in biomedical events, correlating them with the corresponding event triggers.
@@
Compared to flat event structures, texts in the biomedical domain often contain more nested event structures [5]. For example, in benchmark datasets like MLEE [28] and GE11 [16], the nested event structures constitute approximately 25% of all events in the datasets. Figure 1 shows an example of two nested events identified from the given text. The trigger word “promoters” of the Positive Regulation event (E2) serves as the argument role Site for the trigger word “bind”, which triggers the Binding event (E1). Furthermore, “TCF-1 alpha” serves as the argument role Theme for both events (E1 and E2). In addition, a trigger word may also trigger different events in a sentence, depending on its event type or arguments. Events with such structures are called overlapping events. Identifying these complex structures is important for BEE, as it can reveal the underlying relationships between events. For example, a Site2 argument in a Binding event is likely to be a trigger for a Positive Regulation event.
Traditionally, BEE has been formulated as a token classification problem [2, 14]. Although recent classification-based models have demonstrated a certain level of performance in BEE [7, 9], most of them can only handle flat event structures and ignore nested events [5] that appear in the given text. As shown in Figure 2, while classification-based models incorporate structural information, they fail to extract two arguments (highlighted in green) due to their inherent limitation of assigning a single class to each entity. In contrast, generative models with structural information successfully predict all triggers and arguments in the given text. Additionally, many current classification-based methods extract events in a pipeline manner [34], where trigger detection and argument extraction are treated as separate phases, which hinders the learning of shared knowledge between subtasks [33].
@@
Recently, generation-based models have been proposed for event extraction (EE) on general domains [19, 25]. These models typically cast the event extraction task as a sequence generation problem using encoder-decoder based pre-trained language models (PLMs) [29, 17] to output conditional generation sequences. These models often benefit from natural language prompts to learn additional semantic information [8]. Apart from employing natural language prompts to enhance semantic learning from the given context, some recent works such as GTEE [23] and AMPERE [10] have also proposed injecting continuous prompts [21, 11] into their frameworks for better contextual representation learning for event argument extraction. However, most current generation-based EE models [23, 8, 10] have overlooked the incorporation of the structural information. Moreover, Fei et al. [6] pointed out that the mainstay contextual representations in a generative language model’s encoder tend to weaken other high-level information such as structural information, thereby reducing the effects of such information at the decoder. Furthermore, it is also unclear whether the generative language model contains sufficient domain knowledge to represent complex biomedical events. Therefore, it is a challenging task to incorporate structural information effectively into generative BEE models.
@@
To address the above issues, we propose an event structure-aware Generative Biomedical Event Extraction model (GenBEE), which aims to capture complex event structures effectively from biomedical texts. First, GenBEE constructs type description prompts to encapsulate the semantics of type labels. Next, it distills the knowledge from large language models (LLMs), such as GPT-4 [26], by creating event template prompts, which incorporate information such as dependencies among argument roles, to enhance contextual representation learning. Additionally, GenBEE introduces a structural prefix learning module that generates structure-aware prefixes, which are further enriched with structural prompts to provide structure-aware signals.
Figure 2: Comparison of generative and classification-based models in extracting nested events from biomedical text.
@@
GenBEE serves as a comprehensive BEE framework that simultaneously handles event trigger detection and event argument extraction, effectively leveraging the shared knowledge and dependencies between these two subtasks. The main contributions of this paper are summarized as follows:
@@
2 Related Work
Biomedical Event Extraction. Many current methods for biomedical event extraction adopt a “pipeline” strategy, where trigger detection and argument extraction are treated as separate phases. However, “pipeline" methods may result in propagation of errors. To address this issue, Trieu et al. [33] introduced an end-to-end method called DeepEventMine, which simultaneously identifies triggers and allocates roles to entities, thereby mitigating error propagation. However, DeepEventMine requires the full annotation of all entities, which may not always be accessible in datasets. Ramponi et al. [30] considered biomedical event extraction as a sequence labeling task that involves joint training models for trigger detection and argument extraction via multi-task learning. Aside from these classification-based methods, some recent works have formulated event extraction as a question-answering task [35, 18]. This new paradigm transforms traditional classification methods into multiple questioning rounds, producing a natural language answer about a trigger or an argument in each round. Current QA-based event extraction methods are primarily focused on formulating separate questions for different events and argument types. As such, these QA methods lack the capability on capturing dependency information between arguments, which could be effective for assigning argument types.
Generative Event Extraction. Recently, another emerging line of work that treats event extraction as a conditional generation problem [27, 13, 20, 8] has evolved. The development of generation-based event extraction models begins with the reformulation of event extraction problems as a generation task. For example, TANL [27] treated event extraction as translation tasks with augmented labels by using a variety of brackets and vertical bar symbols for representation, and tuned the model to predict these augmented labels. TempGen [13] proposed using templates for constructing role-filler tasks for entity extraction, and generating outputs that fill role entities into non-natural templated sequences. To adapt generation-based models to the task of event argument extraction, BART-Gen [20] investigated generative document-level event argument extraction. However, these methods only use naive natural language prompts and have not considered the incorporation of semantic information into the natural language prompts. To better incorporate label semantics into the model, DEGREE [8] proposed a template-based model that incorporates label semantics into the designed prompts to extract event information. More recently, researchers have started to advance this line of work. For example, GTEE [23] used prefixes [21], which are sequences of tunable continuous prompts, to incorporate contextual type information into the generation-based model. AMPERE [10] proposed to equip the generative model with AMR-aware prefixes to embed the AMR graph, thereby providing more contextual information. To the best of our knowledge, there is no work done on integrating event structural features of complex events into generative models for BEE. Therefore, in this paper, we propose a novel generative model called GenBEE that incorporates complex event structural information for BEE.
3 Method
@@
3.1 Event Prompt Construction
Table 1: Example event prompts for the GE11 dataset.
@@ 3.1 Event Prompt Construction
Table 1 shows some example event prompts for the GE11 dataset. For example, the event type description for the Binding event is “Binding involves two or more molecules coming together to form a complex for various biological processes and signaling pathways.". Note that the type description is obtained according to the textual description of each event type provided in the technical report of the corresponding biomedical event extraction dataset (e.g., GE11 [16]). In the event template, the trigger part is “Event trigger {Trigger}", which remains the same for all event types. The argument part is specific to each event type $e_{i}$. Each {Role_<argument>} serves as a placeholder for the corresponding argument role for the event.
Most previous works [23, 10] proposed to use manually designed templates for event extraction. However, it is challenging and time-consuming to define event templates for each event type in the biomedical event datasets, as domain knowledge is required. In this work, we propose to elicit knowledge from Large Language Models (LLM) GPT-4 [26] with designed prompts to generate the event templates for the event types pre-defined in the biomedical event datasets. Figure 3 shows an input example that we have designed for generating the event template for the Binding event using LLMs. The input prompt consists of an “Instruction” that specifies the generation objective for the biomedical event extraction template and “Valid Roles” that outline the target trigger and argument roles, which are pre-defined by the biomedical event dataset. The output of LLMs is an event template that contains richer semantic information for generative event extraction.
@@
3.2 Structural Prefix Learning
In biomedical texts, events are often complex and interrelated. Identifying the latent relationships among entities across different events could provide important clues for capturing complex event structures, thereby enhancing representation learning. Table 2 shows the structural prompts [37] constructed in the proposed model to capture different event structures for the Binding event. The structural prompts aim to provide descriptions of different event structures and then elicit the PLM to identify complex event structures by capturing the latent relationships among entities across different biomedical events. As shown in Table 2, the structural prompts are the same for all event types, except the corresponding event type name specified between the two tags of <T>. In this work, structural prompts are categorized into three groups: General Events, Overlapping Events, and Nested Events. The General Events describe the structural information of common events, guiding the PLM to identify potential interactions between frequently co-occurring events. The Overlapping Events describe the structural information of overlapping events, instructing the PLM to identify trigger words that commonly trigger overlapping events. The Nested Events describe the structural information of nested events, guiding the PLM to identify entities that frequently serve as argument roles in the events while acting as trigger words in others, or vice versa. Overall, the structural prompts are designed to describe the various structural information of biomedical events for generative event extraction.
Table 2: Example structural prompts for the Binding event.
@@ 3.2 Structural Prefix Learning
General Events (S1) What events are frequently co-occurring with <T>Binding<T> events?
Overlapping Events (S2) What entities are frequently serving as triggers in both <T>Binding<T> events and other event types?
Nested Events (S3) What entities are frequently acting in multiple roles, including as roles in <T>Binding<T> events and differently in other events?
(S4) What entities are frequently acting triggers of <T>Binding<T> events while acting as a role in other events, or vice versa?
@@
3.3 Generative Model Training
@@
4 Experiments
Table 3: Statistics of the datasets.
@@ 4 Experiments
4.1 Experimental Setup
Datasets. We have conducted experiments on the following three publicly available benchmark datasets, namely MLEE, GE11, and PHEE, for biomedical event extraction. MLEE [28] is obtained from the BioNLP-09 Shared Task, which concentrates on recognizing molecular events in biomedical texts. It contains 29 event types and 14 argument roles. We use the train/dev/test split given by the data provider. GE11 [16] is sourced from the BioNLP 2011 Shared Task. This dataset focuses on events associated with transcription factors in the domain of human hematopoietic cells. It contains 9 event types and 10 role arguments. We use the train/dev/test split given by the shared task and evaluate the performance based on the development set, as the test set is unannotated and the official tool for evaluation is no longer available. The PHEE dataset [31] is designed for pharmacovigilance and facilitates a comprehensive analysis of the effects of medical treatments. It contains 2 event types and 16 role arguments, with few nested event structures. We use this dataset to assess the performance of our proposed model on flatter datasets. We use the train/dev/test split provided by TextEE [12] and preprocess each dataset accordingly. Table 3 presents the statistics of the datasets.
Table 4: Experimental results (%) for extracting events based on the MLEE, GE11 and PHEE datasets. The best score is highlighted in bold, and the second-best score is underlined.